Burke Enterprises is considering a machine costing $30 billion that will result in initial after-tax cash savings of $3.7 billion at the end of the first year, and these savings will grow at a rate of 2 percent per year for 11 years. After 11 years, the company can sell the parts for $5 billion. Burke has a target debt/equity ratio of 1.2, a beta of 1.79. You estimate that the return on the market is 7.5% and T-bills are currently yielding 2.5%. Burke has two issuances of bonds outstanding. The first has 200,000 bonds trading at 98% of par, with coupons of 5%, face of $1000, and maturity of 5 years. The second has 500,000 bonds trading at par, with coupons of 7.5%, face of $1000, and maturity of 12 years. Kate, the CEO, usually applies an adjustment factor to the discount rate of +2 for such highly innovative projects. Should the company take on the project?
Answer #1

Cost of equity = risk free rate + Beta * (market return - risk free rate)

Cost of equity = 2.5% + 1.79 * (7.5% - 2.5%) = 11.45%

Market value of first bond (trading at 98% to par value) = 200,000 *1000 * 98% = $196 million

Second bond is trading at par value = 500,000 * 1000 *100% = $500 million

Coupon on first bond = 5% and coupon on second bond = 7.5%

Cost of total debt = (5% * 196 + 7.5% * 500) / (196 + 500)

Cost of total debt = 6.80%

Debt / Equity = 1.2

% of debt = 1.2 / (1.2 +1) = 54.55%

% of equity = 1 / (1.2 +1 ) = 45.45%

Assuming tax rate of 35%

Weighted average cost of capital = 11.45% * 45.45% + 6.8% * (1-35%) * 54.55%

Weighted average cost of capital = 7.62%

Additional adjustment factor of +2 by the CEO, cost of capital = 7.62% + 2% = 9.62%

Applying the project cash flow NPV is -$1.4 billion, hence project should not be accepted

Year Cash flow Discount rate at 9.62% Present value of cash flow
0 -30.0 1.000 -30.0
1 3.7 0.912 3.4
2 3.8 0.832 3.1
3 3.8 0.759 2.9
4 3.9 0.693 2.7
5 4.0 0.632 2.5
6 4.1 0.576 2.4
7 4.2 0.526 2.2
8 4.3 0.480 2.0
9 4.3 0.438 1.9
10 4.4 0.399 1.8
11 9.5 0.364 3.5
NPV -1.6

Excel formula:

A B 1 Year Cash flow Discount rate at 9.62% Present value of cash flow 1/(1+9.62%)^A2 1/(1+9.62%)^A3 1/(1+9.62%)^A4 1/(1+9.62